How many grams are needed to make a 6.5 mol/L solution of calcium chloride in 500 ml flask
1
Expert's answer
2020-11-10T14:03:56-0500

Solution:

"\nu(CaCl_2)=V\times C_M=0.5 L\times 6.5 \frac{mol}{L}=3.25\,mol"

"m(CaCl_2)=\nu\times M=3.25\,mol\times111.08\, \frac{g}{mol}=361.01 \,g"


Answer: 361.01 g